### v1

# A Large-Scale Benchmark of Cross-Modal Learning for Histology and Gene Expression in Spatial Transcriptomics

Abstract
Spatial transcriptomics enables simultaneous measurement of gene expression and tissue morphology, offering unprecedented insights into cellular organization and disease mechanisms. However, the field lacks comprehensive benchmarks for evaluating multimodal learning methods that leverage both histology images and gene expression data. Here, we present HESCAPE, a large-scale benchmark for cross-modal contrastive pretraining in spatial transcriptomics, built on a curated pan-organ dataset spanning 6 different gene panels and 54 donors. We systematically evaluated state-of-the-art image and gene expression encoders across multiple pretraining strategies and assessed their effectiveness on two downstream tasks: gene mutation classification and gene expression prediction. Our benchmark demonstrates that gene expression encoders are the primary determinant of strong representational alignment, and that gene models pretrained on spatial transcriptomics data outperform both those trained without spatial data and simple baseline approaches. However, downstream task evaluation reveals a striking contradiction: while contrastive pretraining consistently improves gene mutation classification performance, it degrades direct gene expression prediction compared to baseline encoders trained without cross-modal objectives. We identify batch effects as a key factor that interferes with effective cross-modal alignment. Our findings highlight the critical need for batch-robust multimodal learning approaches in spatial transcriptomics. To accelerate progress in this direction, we release HESCAPE, providing standardized datasets, evaluation protocols, and benchmarking tools for the community, at https://github.com/peng-lab/hescape
1 Introduction
Figure 1: HESCAPE Benchmark: 4 gene expression encoders and 5 image encoders for digital pathology have been fine-tuned with contrastive pretraining, and evaluated in downstream tasks.
Recent years have witnessed remarkable advances in representation learning for biomedicine, enabling novel strategies to integrate multimodal objectives [30, 12, 46]. Advancements include learning joint representations of whole-slide pathology images and clinical reports [29, 45], integration of pathology image data with bulk gene expression profiles [22], and combining x-ray scans with clinical notes [18]. These successes have been accelerated by the availability of large-scale, carefully curated datasets and robust training paradigms that leverage cross-modal information to yield richer, more generalizable features across domains.
The advent of high-throughput spatial omics techniques, including spatial transcriptomics, has opened new avenues for understanding tissue organization at cellular resolution by leveraging joint histological imaging and molecular profiling data [39, 36]. Such techniques have been particularly impactful in cancer pathology research, where routine pathology analysis is paired with whole-transcriptome measurements from the same biospecimen, yielding rich multimodal molecular data with the potential of uncovering new molecular mechanisms and biomarkers [37]. Furthermore, these data can be integrated with legacy data from pathology biobanks.
Unimodal foundation models have pushed the boundaries for representation learning in pathology [6, 48]. Several instances of large-scale foundation models are available and have demonstrated strong performance across a range of challenging benchmarks [59, 33, 25, 32, 4]. While unimodal foundation models in pathology have shown promise [6, 51, 57, 41, 29], fully harnessing spatial transcriptomics demands new multimodal strategies capable of learning from both tissue images and gene expression measurements simultaneously. Such approaches have the potential to highlight principles of tissue architecture that might remain obscured in single-modality analyses, as well as enhance model performance in unimodal predictive tasks.
To elucidate this challenge, we curated a large-scale dataset of over $\sim 720k$ patch-level image–gene pairs. These data span multiple organs and originate from 6 different 10x Xenium in-situ panels, collected across 54 patients and 57 individual samples. Filtering out replicates with different gene panels and single gene panel datasets, we use $\sim 620k$ image-gene pairs for the current benchmark. Through systematic examination of image and gene expression encoder combinations across a comprehensive set of configurations, we present an extensive benchmark of modality-specific encoder performance in contrastive pretraining settings. Our results demonstrate that dataset-specific pretraining of gene expression encoders has the greatest impact on cross-modality retrieval performance. Additionally, we emphasize the critical importance of addressing batch effects in these encoders. For the gene expression modality specifically, the ability to directly model batch effects provides significant performance improvements in cross-modality retrieval tasks.
We directly evaluate the top-performing methods from contrastive pretraining cross-modality retrieval tasks in real-world downstream applications. Specifically, we assess how contrastive pretraining on spatial transcriptomics datasets impacts performance on two downstream tasks: mutation-specific tumor classification and gene expression prediction from morphology Fig.˜1.
Building on these findings, we propose HESCAPE (H&E + Spatial ContrAstive PrEtraining), a framework for contrastive pretraining on spatial transcriptomics data. HESCAPE provides a curated dataset with one-to-one mappings between histological patches and gene expression vectors, along with standardized performance metrics for evaluating new architectures. By making this open benchmark available on Hugging Face, we aim to accelerate the development of robust, clinically applicable spatial omics models that integrate seamlessly with widely adopted imaging encoders. This resource represents a step toward enabling the computer vision community to tackle domain-specific challenges in spatial transcriptomics, thereby advancing both fundamental multi-omics research and translational applications in oncology and beyond.
2 Related work
2.1 Image encoders
The success of large-scale self-supervised learning combined with the availability of large numbers of unlabeled image patches from histology slides initiated the development of domain-specific feature encoders. These models are trained on millions of image patches from hundreds of thousands of patients using contrastive methods such as MoCov3 (RetCCL [52], CTransPath [51]) or knowledge distillation approaches such as DINO [5] or DINOv2 [35] (Phikon, UNI [6], Virchow [48], H-Optimus-0 [41, 13], Prov-GigaPath [57], Atlas [1]). These pathology-specific encoders provide better and more robust image representations, and enhance critical downstream tasks, such as slide-level biomarker prediction [49] or clinical report generation [45].
2.2 Gene encoders
Single-cell transcriptomics consists of gene expression measurements at single-cell resolution. The data is typically sparse and contains signals reflecting both biological variation (e.g., tissue and cell state variation) and technical variation (e.g., sample preparation, sequencing depth). Variational autoencoders (VAEs) based on multilayer perceptrons (MLPs) such as scVI [28] or DRVI [34] learn a low-dimensional embedding by aiming to reconstruct the gene expression at input. Recently, researchers have developed foundation models for transcriptomics (Geneformer [10], scGPT [8], scFoundation [14], UCE [40], Nicheformer [42], scGPT-spatial [50]) that encode information from tens of millions of cells. While these models show that large-scale pretraining can provide benefit in downstream tasks in single-cell genomics, despite limited model capacity [11], their ability to encode molecular variation in a multimodal representation learning setting has not been tested to date.
2.3 Multi-modal models in histopathology
Combining pre-trained text encoders, image encoders, and gene encoders offers the opportunity to learn richer representations of multimodal patient data such as pathology reports, pathology images, and bulk RNA sequences. Recent models have proposed to align pathology images with text on patch-level (PLIP [17], QUILT-1M [19], CONCH [29], MUSK [55]) or on slide-level (PRISM [44], GigaPath [57], TITAN [12]). Beyond morphological descriptions and clinical reports, pathology slides can be aligned with bulk gene expression (TANGLE [22], FSM [61], THREADS [46], SURVPATH [23]), or immunohistochemistry (MADELEINE [24]). To take a step further, multi-modal integration with 3D spatial-omics has shown promising results (VORTEX [2]).
2.4 Multi-modal in spatial-omics data
Several models have been proposed for integrating the dual modality of spatial omics data [36, 39]. A key challenge is the prediction of gene expression profiles from morphology [15, 43, 3, 60]. This task was recently proposed as a foundational benchmark for multimodal histology and gene expression integration [21]. Other studies suggest pretraining the multimodal representation with contrastive learning [26, 56, 7], showing improved performance for the gene expression prediction task. Nevertheless, a large-scale evaluation of multimodal pretraining in spatial omics, which leverages different image and gene expression encoders as well as a large and diverse dataset across various transcriptomic panels, has not been reported. Additionally, the importance of performing multimodal pretraining on spatial-omics data, the correct selection of encoders and its corresponding impact on unimodal downstream task performance for histopathology is yet to be studied.
3 Large-scale Contrastive Pretraining Benchmark For Spatial Transcriptomics
In this work, we devised a large-scale benchmark to test the effectiveness of contrastive pretraining methods for learning multimodal representations of spatial transcriptomics data. In the following section, we describe the datasets, image encoder, and gene expression decoders employed in the benchmark.
3.1 Dataset
We curated a large collection of spatial transcriptomics data from 10x Genomics Xenium consisting of solely human tissues. The HESCAPE collection has 6 datasets, each identified by a subset of samples that share the same gene panels used to measure the samples. To highlight batch effects both across organs and collection sites, as well as to evaluate their effect on multi-modal learning, 3 datasets consist of organ-specific Xenium samples and 3 datasets consist of pan-organ Xenium samples. These samples have been pooled from different sites such as HEST benchmark [21] and other publicly available sources. It consists of a total of 729962 patches (623623 without duplicates). A full description of the dataset curation and preprocessing can be found in Tab.˜1 and Supplementary Sec.˜7.1.
Table 1: HESCAPE Dataset characteristics across 10x Xenium gene panels showing number of genes, patches, tissue sections, and samples/donors.
Panel Genes Patches Sections Donors
5K 5001 178817 7 7
Breast 280 127960 5 5
ImmOnc 380 67050 5 5
Lung 343 56689 20 19
Colon 322 61067 5 3
Multi 377 132040 15 15
3.2 Models
We evaluate diverse combinations of image and gene encoder models using contrastive learning across all 6 datasets. To support future development, we open-source the benchmarking code and provide detailed instructions, fostering collaboration and the adoption of innovative, batch-robust gene encoder practices. The HESCAPE benchmark is designed as a plug-and-play framework, enabling seamless integration and rapid comparison of novel encoders from either modality.
3.2.1 Image encoders
We include five state-of-the-art image encoder models, each trained on a minimum of 15 million pathology patches collected from diverse sources.
CTransPath is a Swin Transformer trained on 15 million patches from 30,000 whole slide images (WSIs) in The Cancer Genome Atlas (TCGA) and Pathology Artificial Intelligence Platform (PAIP) public databases [51].
UNI is a vision transformer (ViT) trained on 100 million patches from more than 100,000 H&E stained WSIs spanning 20 major tissue types using the DINOv2 self-supervised framework based on knowledge distillation [6].
CONCH is a ViT trained using vision-language alignment with CLIP on 1.15 million image-caption pairs curated from PubMed and medical textbooks [29].
Gigapath is a ViT model trained on approximately 1.1 billion pathology patches derived from 170,000 WSIs. Similar to UNI, its patch encoder is trained using the DINOv2 self-supervision framework [57].
H0mini is an 86 million parameters ViT, distilled from 1.1 billion parameters original H-Optimus-0 [41], trained on a proprietary dataset of over 500,000 H&E stained WSIs [13].
3.2.2 Gene encoders
Despite the limited number of dedicated gene encoders for spatial technologies, we evaluate four models that are straightforward to train or fine-tune on spatial datasets. These models have either been developed as foundational models for single-cell or spatial genomics, or they have demonstrated robustness in extracting meaningful embeddings from gene expression data, whether at the single-cell or spatial level.
MLP: A multilayer perceptron (MLP) is a standard baseline for gene expression data. Our implementation uses a simple two-layer MLP with a hidden layer twice the size of the input dimension. This model is trained from scratch alongside the contrastive training on the specific gene expression dataset of interest, providing a straightforward benchmark for gene coding performance.
DRVI: This is an unsupervised VAE-based generative model that learns nonlinear, disentangled representations of single-cell data, designed as an extension to scVI [34]. We pretrain DRVI independently on each gene expression training dataset separately.
scFoundation: A transformer-based foundation model trained on a large collection of single-cell data. Since scFoundation has shown its generalizable feature representations extend to other modalities like bulk RNA-seq data, we apply the scFoundation model directly without additional spatial fine-tuning [14].
Nicheformer: A transformer-based model based on BERT and adapted to both dissociated single-cell and targeted spatial transcriptomics data (e.g., 10x Xenium), Nicheformer has demonstrated exceptional zero-shot and fine-tuning performance in spatially relevant downstream tasks [42]. We use the frozen nicheformer model in zero-shot settings for our benchmark.
3.3 Pretraining
Our experimental setup is based on a contrastive multi-modal pretraining framework. Each input to the vision encoder is an image patch $I_{i}$ surrounding a Xenium patch, and the encoder outputs a feature vector $\mathbf{v}_{i}\in\mathbb{R}^{d}$, where $d$ is the feature dimension. Each gene encoder processes the corresponding paired expression $G_{i}$, with pre-processing steps such as normalization or ranking applied when necessary (based on the encoder requirements). It yields embeddings $\mathbf{g}_{i}\in\mathbb{R}^{d}$, where $d$ is the same feature dimension as for the image encoders.
3.3.1 Losses
We leverage the foundational CLIP (Contrastive Language-Image Pre-Training) [38] objective, as well as the newer SigLip (Sigmoid Loss for Language-Image Pre-training) [58] objective.
Specifically, for a minibatch of images and gene expression pairs =$\mathcal{B}=\left\{\left(I_{1},G_{1}\right),\left(I_{2},G_{2}\right),\ldots\right\}$, the CLIP learning objective for the image-gene expression pair v2g is:
$L_{\texttt{CLIP}^{\texttt{v2g}}}=-\frac{1}{|\mathcal{B}|}\sum_{i=1}^{|\mathcal{B}|}\log\frac{\exp(\tau\langle\mathbf{v}_{i},\mathbf{g}_{i}\rangle)}{\sum_{j=1}^{|\mathcal{B}|}\exp(\tau\langle\mathbf{v}_{i},\mathbf{g}_{j}\rangle)}$ (1)
where, again, $\mathbf{v}_{i}\in\mathbb{R}^{d}$ and $\mathbf{g}_{i}\in\mathbb{R}^{d}$ denote the image embeddings and gene embeddings from their respective encoders. The temperature $\tau$ is a learnable parameter.
For both losses, the global loss is
$L_{\texttt{CLIP}}=\frac{1}{2}\left[L_{\texttt{v2g}}+L_{\texttt{g2v}}\right]$ (2)
where the inner product in $L_{\texttt{g2v}}$ is $\langle\mathbf{g}_{i},\mathbf{v}_{j}\rangle$. In this work, we utilize both losses across the pretraining experiments to understand their performance in our data and batch size regime (Supplementary Tab.˜10). We leverage the openclip implementations in Ilharco et al. [20] for multi-GPU loss computation of CLIP and SigLip.
In the pretraining phase, we experiment with multimodal encoders both frozen and fine-tuned, where we do full fine-tuning for the MLP and DRVI and parameter-efficient fine tuning (PEFT) using LoRA (Low-Rank Adaptation)[16] for all image models, Nicheformer, and scFoundation. Additional description of the experimental configurations are shown in  Sec.˜8.2.3. For the final benchmark shown in Tab.˜2, we keep all gene encoders frozen except for the MLP baseline, which is trained from scratch alongside the contrastive pretraining.
3.3.2 Metrics
[…]
3.4 Downstream evaluations
3.4.1 Gene Mutation Prediction
Driver mutations are a hallmark of cancer. These genetic alterations directly promote the development and growth of cancer cells. They are often associated with tissue morphology (that is, the histological phenotype) and can be identified using deep learning models [49]. In cancer diagnostics, this could eliminate costly and time-consuming molecular tests or genetic assays. For instance, in colorectal cancer (CRC), a mutation in BRAF is associated with a poorer prognosis. A vision model that can reliably detect BRAF mutations in routine histology images could identify high-risk patients earlier. This would guide the selection of targeted therapies and ultimately improve treatment outcomes. Here, we analyze whether pathology foundation models trained on image data alone are sufficient for biomarker detection and if they can be further improved by aligning gene expression data.
We benchmark multimodal alignment on nine biomarker detection tasks, three biomarkers for CRC: microsatellite instability (MSI), BRAF, and KRAS status; three biomarkers for breast adenocarcinoma (BRCA): ER, PR, and HER2 status; and three biomarkers for lung adenocarcinoma (LUAD): EGFR, KRAS, and TP53 status. All tasks are tested on H&E-stained FFPE WSIs from the respective cohorts of TCGA [53, 27].
3.4.2 HEST Benchmark
Fine-tuned image encoders are evaluated on the HEST benchmark [21] by regressing patch-level gene expression from image patches. The benchmark comprises nine tasks that span various tissue types and cancer subtypes from both 10x xenium and visium datasets. Each task predicts the expression of the top 50 most variable genes from 224$\times$224 pixel histology patches 112 $\mu$m $\times$ 112 $\mu$m at 20x magnification. The features extracted at the patch level are then passed to a ridge regression model for gene expression prediction.
To ensure a fair comparison and avoid data leakage, we maintained the identical train/test splits used in the pretraining phase, which vary slightly from the HEST benchmark. Further descriptions on dataset creation and train/test splits have been reported in the Supplementary Sec.˜7.1.
Here, we assess whether HESCAPE-aligned image encoders trained with spatial transcriptomics yield better expression prediction performance. We report Pearson’s correlation coefficient (PCC) and mean squared error (MSE) between predicted and measured gene expression, averaged across genes and cross-validation folds (patient-level splits).
4 Results
4.1 Contrastive Pretraining
Our large-scale pretraining benchmark demonstrates that the gene encoder DRVI emerges as the key determinant of performance improvement across all tissue panels and both image-to-gene (I2G) and gene-to-image (G2I) metrics (Tab.˜2 and Supplementary Tab.˜7-Tab.˜8). Notably, DRVI paired with the image encoders Gigapath, H0mini, and UNI consistently achieves the top-performing Recall@5 scores across datasets. This superior performance is likely attributable to DRVI’s pretraining on the gene expression modality of the training set and its capability to handle batch effects effectively.
When examining pre-trained gene foundation models, our findings reveal that despite their pretraining on large-scale datasets, they cannot surpass VAE-based models like DRVI pretrained on the dataset of interest. However, despite the training of MLP baselines on the dataset of interest, foundation models do demonstrate substantial improvements over them, with performance gains ranging from approximately 2-5x across different tissue types. Among the foundation models in our benchmark, Nicheformer consistently outperforms scFoundation across all evaluated datasets, highlighting the advantage of pretraining on spatial transcriptomics data compared to models trained exclusively on single-cell genomics data.
The results also reveal significant heterogeneity in performance across tissue types, with lung tissue showing the highest overall performance (Recall@5 up to 0.709 for Gigapath-DRVI G2I) and the 5K dataset showing the most challenging retrieval task. Interestingly, while the I2G and G2I tasks generally show comparable performance patterns, some tissue-specific variations emerge, particularly in breast and lung datasets, where certain models show slightly asymmetric performance between the two retrieval directions.
Additionally, we perform several ablation studies to observe contributing factors for optimal pretraining. From these ablations, we observe, an optimal combination of MLP projection, CLIP loss with fine-tunable image and gene encoders performs well across all spatial datasets. However, limited computational resources require us to perform locked gene-image tuning to improve the image encoders. The ablations have been reported in Sec.˜8.2 and Supplementary Tab.˜9, Tab.˜10, and Tab.˜11.
Table 2: Complete Test Recall@5 Results for both Image-to-Gene (I2G) and Gene-to-Image (G2I) tasks across different tissue panels. Experiments with "—" indicate out of memory issues during training. Best results are in bold, second-best are underlined.
5K Multi ImmOnc Colon Breast Lung
Model I2G G2I I2G G2I I2G G2I I2G G2I I2G G2I I2G G2I
MLP-CTransPath 0.103 0.106 0.138 0.098 0.110 0.094 0.098 0.122 0.116 0.117 0.103 0.126
MLP-CONCH 0.228 0.228 0.241 0.178 0.187 0.130 0.300 0.258 0.383 0.260 0.443 0.418
MLP-Gigapath 0.257 0.257 0.297 0.215 0.179 0.132 0.313 0.297 0.390 0.288 0.510 0.493
MLP-H0mini 0.235 0.235 0.209 0.153 0.173 0.119 0.296 0.291 0.309 0.235 0.358 0.336
MLP-UNI 0.247 0.246 0.255 0.171 0.243 0.130 0.320 0.317 0.346 0.248 0.493 0.493
scFoundation-CTransPath 0.119 0.126 0.105 0.098 0.138 0.122 0.125 0.121
scFoundation-CONCH 0.219 0.177 0.287 0.262 0.331 0.309 0.503 0.458
scFoundation-Gigapath 0.251 0.207 0.294 0.249 0.348 0.365 0.590 0.543
scFoundation-H0mini 0.206 0.171 0.315 0.272 0.388 0.377 0.427 0.345
scFoundation-UNI 0.236 0.173 0.297 0.244 0.358 0.351 0.543 0.479
Nicheformer-CTransPath 0.105 0.115 0.126 0.117 0.127 0.127 0.092 0.106 0.110 0.126 0.122 0.138
Nicheformer-CONCH 0.237 0.274 0.258 0.286 0.224 0.247 0.267 0.253 0.366 0.410 0.412 0.496
Nicheformer-Gigapath 0.241 0.255 0.274 0.285 0.247 0.267 0.261 0.269 0.414 0.447 0.473 0.554
Nicheformer-H0mini 0.243 0.273 0.261 0.277 0.212 0.215 0.290 0.278 0.418 0.451 0.424 0.498
Nicheformer-UNI 0.259 0.291 0.269 0.288 0.243 0.261 0.252 0.250 0.416 0.442 0.449 0.538
DRVI-CTransPath 0.106 0.116 0.116 0.147 0.138 0.135 0.111 0.123 0.162 0.144 0.143 0.163
DRVI-CONCH 0.266 0.316 0.298 0.363 0.300 0.290 0.356 0.362 0.397 0.397 0.539 0.597
DRVI-Gigapath 0.315 0.359 0.322 0.417 0.344 0.334 0.388 0.394 0.461 0.436 0.649 0.709
DRVI-H0mini 0.299 0.321 0.271 0.342 0.287 0.267 0.412 0.397 0.465 0.461 0.562 0.612
DRVI-UNI 0.322 0.341 0.312 0.396 0.326 0.318 0.404 0.401 0.450 0.436 0.610 0.678
4.2 Gene Mutation Prediction
We tested whether contrastive pretraining with gene expression profiles provides improvements on whole-slide image (WSI) level classification tasks. To test how gene-panel specific pretraining affects the downstream performance across various organs, we choose our top 2 performing models Gigapath-DRVI and UNI-DRVI trained on the Human Multi-Tissue Panel. Additionally, to show how gene-encoder selection can affect downstream tasks, for the Human 5K Panel, we selected Gigapath-DRVI as our top model and UNI-Nicheformer as our sub-optimal model. The Human 5K Panel and the Human Multi-Tissue Panel were used as the common dataset for training to show that performance gains in gene mutation prediction task are not specific to certain organ datasets. We conducted downstream evaluations on mutation prediction for three biomarkers for CRC, three for BRCA, and three for LUAD (Sec.˜3.4.1).
The results in Tab.˜3 reveal a nuanced pattern regarding the benefits of contrastive pretraining with gene expression encoders. For MSI prediction, Gigapath-DRVI achieves the best performance in F1 score (0.558), outperforming the baseline Gigapath model (0.484), representing a 15.3% improvement. Similarly, for BRAF prediction, UNI-DRVI (0.395) outperforms the respective baseline model (0.325) by 21.5% and the best baseline model Gigapath (0.381) by 2.6%. However, for KRAS prediction in CRC, both baseline models, UNI (0.495) and Gigapath (0.491), outperform their gene expression-pretrained counterparts, Gigapath-DRVI (0.474) and UNI-DRVI (0.364).
Table 3: Performance comparison of different encoders on CRC classification tasks. Results show F1 scores (mean(std)) for MSI, BRAF, and KRAS prediction tasks. Best results are in bold, second-best are underlined.
Encoder MSI BRAF KRAS
Gigapath 0.484(0.059) 0.381(0.108) 0.491(0.106)
UNI 0.472(0.099) 0.325(0.137) 0.495(0.088)
Human Multi-Tissue Panel
Gigapath-DRVI 0.558(0.091) 0.394(0.104) 0.474(0.097)
UNI-DRVI 0.465(0.098) 0.395(0.076) 0.364(0.210)
Human 5K Panel
Gigapath-DRVI 0.473(0.093) 0.353(0.041) 0.421(0.098)
UNI-Nicheformer 0.416(0.032) 0.338(0.092) 0.419(0.156)
Similarly, biomarker prediction in BRCA demonstrates a comparable pattern of performance (Tab.˜4). For ER status prediction, Gigapath-DRVI achieves the highest F1 score of 0.898, representing a slight improvement over the baseline Gigapath (0.890), consistent with the UNI variants. For PR status prediction, Gigapath-DRVI again shows the best performance with an F1 score of 0.829, outperforming both baseline models, Gigapath (0.811) and UNI (0.807). However, for HER2 status prediction, the baseline models demonstrate superior performance, with UNI achieving the highest F1 score of 0.439, compared to the gene expression-pretrained models, Gigapath-DRVI (0.401) and UNI-DRVI (0.368).
Table 4: Performance comparison of different encoders on BRCA classification tasks. Results show F1 scores (mean(std)) for ER, PR, and HER2 prediction tasks. Best results are in bold, second-best are underlined.
Encoder ER PR HER2
Gigapath 0.890(0.022) 0.811(0.019) 0.388(0.053)
UNI 0.857(0.017) 0.807(0.038) 0.439(0.021)
Human Multi-Tissue Panel
Gigapath-DRVI 0.898(0.028) 0.829(0.039) 0.401(0.042)
UNI-DRVI 0.887(0.043) 0.817(0.022) 0.368(0.058)
Human 5K Panel
Gigapath-DRVI 0.872(0.018) 0.799(0.026) 0.407(0.023)
UNI-Nicheformer 0.858(0.022) 0.808(0.030) 0.433(0.023)
Finally, WSI gene mutation prediction in LUAD results in the same biomarker-specific pattern, though with more pronounced improvements depending on the vision backbone (Tab.˜5). For two out of three biomarkers, pretraining with gene expression encoders demonstrates substantial performance gains. Most notably, for EGFR prediction, Gigapath-DRVI achieves an F1 score of 0.299, representing a dramatic 100% improvement over the baseline Gigapath (0.149). Similarly, for KRAS prediction, Gigapath-DRVI from the 5K panel achieves the best performance with an F1 score of 0.433, outperforming the baseline Gigapath (0.367) by 18%. However, for TP53 prediction, there isn’t a substantial improvement of the trained models over the baseline Gigapath which has an F1 score of 0.705, while both gene expression-pretrained variants score similarly with high variance across different training folds, Gigapath-DRVI (0.695) and UNI-DRVI (0.649). This pattern could be observed due to variations in self-supervised training data used for UNI and Gigapath baselines. The pattern also reinforces the biomarker-specific nature of the benefits observed across all three cancer types.
It should be noted that in the case of LUAD WSI gene mutation prediction, we only had one sample of LUAD measured with Xenium technologies for each of the two gene panels under consideration (Human 5K Panel and Human Multi-Tissue Panel). Hence, this represents pretraining in a pan-cancer setting rather than lung-specific pretraining, which may explain some of the performance variations observed compared to the more tissue-specific pretraining scenarios in CRC and BRCA analyses.
Table 5: Performance comparison of different encoders on LUAD classification tasks. Results show F1 scores (mean(std)) for EGFR, KRAS, and TP53 prediction tasks. Best results are in bold, second-best are underlined.
Encoder EGFR KRAS TP53
Gigapath 0.149(0.0797) 0.367(0.1437) 0.705(0.0567)
UNI 0.349(0.070) 0.382(0.095) 0.693(0.057)
Human Multi-Tissue Panel
Gigapath-DRVI 0.299(0.0432) 0.371(0.0970) 0.668(0.0662)
UNI-DRVI 0.329(0.0610) 0.379(0.0839) 0.649(0.0769)
Human 5K Panel
Gigapath-DRVI 0.328(0.0383) 0.433(0.0791) 0.695(0.0656)
UNI-Nicheformer 0.301(0.0437) 0.387(0.1156) 0.658(0.0581)
An important limitation of this analysis is that the expression profiles of the target genes associated with the genomic abnormalities were not available in our Xenium datasets used for pretraining. For instance, when predicting TP53 mutations, the actual TP53 gene expression values were not measured in the spatial transcriptomics samples. Consequently, the models could only leverage latent correlation structures between the measured genes in the spatial transcriptomics data and the genomic abnormalities of interest, rather than direct gene-to-mutation associations.
This constraint may explain some of the observed variability in predictive performance across different biomarkers. Mutations in genes with stronger co-expression networks or clearer downstream effects on the measured gene panel are more likely to be predicted accurately through correlated expression patterns. In contrast, mutations in genes with weaker or more complex regulatory relationships to the measured genes might be harder to detect through proxy gene expression patterns. For instance, in BRCA, ESR1 gene is tightly linked to the ER receptor status, PGR to PR status and ERBB2 to HER2 receptor status. All three of these genes are included in the Multi-tissue and Prime 5K gene panels.
In summary, this analysis demonstrates that while integrating molecular knowledge into vision models can improve WSI mutation prediction for certain biomarkers, the effectiveness is highly biomarker-specific. The success appears to depend on both the strength of underlying biological relationships between histological features and gene expression variation, as well as the presence of measurable proxy signals within the available gene expression data that correlate with the target genomic abnormalities.
4.3 Gene Expression Prediction
Next, we investigated whether contrastive pretraining improves the downstream task of predicting gene expression directly from histological images. Based on the superior retrieval performance demonstrated in Tab.˜2, we selected the top-performing pretrained multimodal encoders and evaluated their image encoders on gene expression prediction using a ridge regression head as performed in the HEST gene expression prediction benchmark [21]. To ensure a fair comparison and avoid data leakage, we maintained the identical train/test splits used in the pretraining phase. The gene expression prediction performance, evaluated using Pearson correlation coefficient (PCC) and mean squared error (MSE), is reported in Tab.˜6.
Table 6: Gene Expression Prediction Task as given by HEST benchmark. Best results are in bold.
5K ImmOnc Breast Lung
Model PCC MSE PCC MSE PCC MSE PCC MSE
H0mini 0.321(0.022) 2.726(0.831) 0.571(0.026) 6.198(0.137) 0.639(0.050) 15.237(3.962) 0.628(0.013) 10.466(0.109)
Gigapath 0.338(0.021) 2.408(0.877) 0.546(0.030) 6.167(0.210) 0.607(0.049) 15.753(5.160) 0.609(0.024) 10.604(0.406)
UNI 0.319(0.031) 2.535(1.164) 0.546(0.020) 5.984(0.163) 0.605(0.023) 14.428(5.404) 0.607(0.019) 10.705(0.593)
Human Multi-Tissue Panel
DRVI-H0mini 0.247(0.048) 2.713(0.626) 0.476(0.038) 6.204(0.212) 0.520(0.002) 10.849(3.173) 0.479(0.005) 11.152(1.008)
DRVI-Gigapath 0.277(0.020) 2.461(0.712) 0.479(0.028) 6.011(0.075) 0.560(0.026) 10.600(3.593) 0.549(0.010) 10.984(1.182)
DRVI-UNI 0.231(0.062) 3.126(0.499) 0.487(0.028) 5.986(0.134) 0.568(0.048) 10.362(3.661) 0.548(0.007) 10.467(0.217)
Human 5K Panel
DRVI-H0mini 0.218(0.032) 2.895(0.920) 0.493(0.038) 5.974(0.391) 0.519(0.039) 11.260(3.737) 0.501(0.005) 11.339(0.757)
DRVI-Gigapath 0.232(0.018) 2.665(0.884) 0.530(0.039) 6.230(0.219) 0.571(0.062) 11.228(4.492) 0.563(0.009) 10.485(1.225)
DRVI-UNI 0.264(0.035) 2.661(0.671) 0.495(0.023) 6.280(0.269) 0.592(0.065) 11.904(5.408) 0.543(0.000) 10.845(1.146)
Surprisingly, the results in  Tab.˜6 reveal a counterintuitive pattern: contrastive pretraining with gene expression encoders does not consistently improve direct gene expression prediction from histological images. Across all four datasets, the baseline image encoders (without gene expression pretraining) generally achieve superior or comparable performance compared to their pretrained counterparts. For instance, in the 5K dataset, the baseline Gigapath model achieves the highest PCC of both image encoders pre-trained on the Multi and 5K variants. This pattern is consistent across other datasets, where the baseline H0mini achieves the best performance in the ImmOnc, Breast, and Lung datasets.
This negative result is counterintuitive, given that cross-modal pretraining should, in principle, improve the prediction of one modality from the other. We propose two possible explanations for this discrepancy from our initial hypothesis.
First, the strong batch effects present in the gene expression modality may skew the image encoder’s learned representations, causing it to prioritize batch-specific patterns over generalizable biological features, thereby reducing performance on the test set. Second, we hypothesize that the alignment process itself may be detrimental to the image encoder’s representational quality. During contrastive pretraining, the image encoder may be forced to discard rich morphological and spatial information that is crucial for gene expression prediction but not necessary for the contrastive task.
To support this hypothesis, we present an analysis on evaluating the magnitude of batch effects in Figure 2 and Supplementary Sec.˜8.3, leveraging the Silhouette Batch score from scib [31].
Figure 2: Silhouette Batch score against Mean Recall@5 for Gigapath-DRVI encoder models. Interestingly, there is a clear trend between the cross-modal retrieval task and the magnitude of batch effect in the train/val/test split.
Datasets with better batch integration (higher Silhouette Batch scores) consistently achieve superior cross-modal retrieval performance, suggesting that technical artifacts in the gene expression modality can substantially hinder the learning of meaningful cross-modal representations.
This potentially creates a trade-off between representation alignment and representation quality: while the pretrained encoder learns features that are better aligned with gene expression patterns, these aligned features may be less informative for the downstream prediction task compared to the richer, more generalizable features learned by the baseline encoder. In summary, the biological correlation between histological features and gene expression may be weaker than initially assumed or may be strongly affected by experimental noise such as batch effects. This creates a challenging environment for cross-modal pretraining, where strong domain-specific effects in each modality can interfere with learning meaningful cross-modal relationships. Our findings suggest that future work should focus on developing multimodal encoders that explicitly account for and mitigate domain-specific effects within each modality. Such approaches could potentially achieve greater robustness to batch effects while still leveraging the benefits of contrastive pretraining with spatial transcriptomics data. This would involve disentangling biological signal from technical noise either in an additional pretraining stage before cross-modal alignment, or online during contrastive pretraining [9], potentially leading to more effective and generalizable multimodal representations.
5 Conclusion
In this work, we conducted a large-scale benchmark of contrastive pretraining encoders for cross-modal alignment of morphology and gene expression modalities in histopathology. We demonstrated improved performance on cross-modal retrieval tasks when encoders were pretrained on the dataset of interest and capable of handling batch effects appropriately. Among the foundation models tested, those directly pretrained on spatial transcriptomics data showed superior performance, highlighting the importance of including spatial transcriptomics data in the pretraining corpus for foundational single-cell genomics representations.
Furthermore, we evaluated pretrained image encoders on two downstream tasks: gene mutation classification and gene expression prediction for specific biomarkers. For mutation classification, cross-modally pretrained image encoders showed convincing improvements in performance. However, for gene expression prediction, the results were mixed, with baseline encoders often outperforming their cross-modally pretrained counterparts. This finding conflicts with findings in the literature where pretraining with gene expression appears to always improve image-to-gene expression prediction tasks [7, 56, 26].
We hypothesize that these results are driven by strong domain-specific effects in the gene expression modality, particularly batch effects that may interfere with the learning of generalizable cross-modal representations. Our analysis provides evidence for such effects and suggests that further investigation is required to directly address domain-specific effects towards learning robust and transferable multimodal representations of morphology and gene expression. Such advances would have a significant translational impact in digital pathology and patient stratification applications.
6 Acknowledgments
R.H.G., G.P, S.J.W., M.T were supported by the Helmholtz Association under the joint research school “Munich School for Data Science - MUDS”. S.J.W. was supported by Add-on Fellowship of the Joachim Herz Foundation. The compute resources for the project were supported by de.NBI Cloud within the German Network for Bioinformatics Infrastructure (de.NBI), ELIXIR-DE (Forschungszentrum Jülich) and Helmholtz Association’s Initiative and Networking Fund on the HAICORE@FZJ partition.

### v2

@@
 Figure 1: HESCAPE Benchmark: 4 gene expression encoders and 5 image encoders for digital pathology have been fine-tuned with contrastive pretraining, and evaluated in downstream tasks.
-Recent years have witnessed remarkable advances in representation learning for biomedicine, enabling novel strategies to integrate multimodal objectives [30, 12, 46].
-Advancements include learning joint representations of whole-slide pathology images and clinical reports [29, 45], integration of pathology image data with bulk gene expression profiles [22], and combining x-ray scans with clinical notes [18].
-These successes have been accelerated by the availability of large-scale, carefully curated datasets and robust training paradigms that leverage cross-modal information to yield richer, more generalizable features across domains.
+Recent advances in biomedical representation learning have enabled joint modeling of pathology images with reports [29, 46], bulk gene expression [22], and x-ray scans with clinical notes [18], largely driven by large-scale curated datasets and robust training paradigms [12, 30, 47].
 The advent of high-throughput spatial omics techniques, including spatial transcriptomics, has opened new avenues for understanding tissue organization at cellular resolution by leveraging joint histological imaging and molecular profiling data [40, 37].
@@
 Furthermore, these data can be integrated with legacy data from pathology biobanks.
-Unimodal foundation models have pushed the boundaries for representation learning in pathology [6, 48].
-Several instances of large-scale foundation models are available and have demonstrated strong performance across a range of challenging benchmarks [59, 33, 25, 32, 4].
-While unimodal foundation models in pathology have shown promise [6, 51, 57, 41, 29], fully harnessing spatial transcriptomics demands new multimodal strategies capable of learning from both tissue images and gene expression measurements simultaneously.
-Such approaches have the potential to highlight principles of tissue architecture that might remain obscured in single-modality analyses, as well as enhance model performance in unimodal predictive tasks.
-To elucidate this challenge, we curated a large-scale dataset of over $\sim 720k$ patch-level image–gene pairs.
-These data span multiple organs and originate from 6 different 10x Xenium in-situ panels, collected across 54 patients and 57 individual samples.
-Filtering out replicates with different gene panels and single gene panel datasets, we use $\sim 620k$ image-gene pairs for the current benchmark.
-Through systematic examination of image and gene expression encoder combinations across a comprehensive set of configurations, we present an extensive benchmark of modality-specific encoder performance in contrastive pretraining settings.
+Pathology foundation models have achieved strong performance across benchmarks [6, 49, 60, 34, 25, 32, 4].
+To fully harness the information provided by spatial transcriptomics, we require novel multimodal strategies capable of learning of integrating tissue images with gene expression.
+To elucidate this challenge, we curated a large-scale dataset of over $\sim 620k$ patch-level image–gene pairs across multiple organs originating from 6 different 10x Xenium in-situ panels, collected across 54 patients and 57 individual samples.
+This dataset enable us to perform a large-scale benchmark of image and gene encoders using contrastive pretraining.
 Our results demonstrate that dataset-specific pretraining of gene expression encoders has the greatest impact on cross-modality retrieval performance.
@@
 We leverage the foundational CLIP (Contrastive Language-Image Pre-Training) [39] objective, as well as the newer SigLip (Sigmoid Loss for Language-Image Pre-training) [59] objective.
-Specifically, for a minibatch of images and gene expression pairs =$\mathcal{B}=\left\{\left(I_{1},G_{1}\right),\left(I_{2},G_{2}\right),\ldots\right\}$, the CLIP learning objective for the image-gene expression pair v2g is:
+Specifically, for a minibatch of images and gene expression pairs $\mathcal{B}=\left\{\left(I_{1},G_{1}\right),\left(I_{2},G_{2}\right),\ldots\right\}$, the CLIP learning objective for the image-gene expression pair v2g is:
 $L_{\texttt{CLIP}^{\texttt{v2g}}}=-\frac{1}{|\mathcal{B}|}\sum_{i=1}^{|\mathcal{B}|}\log\frac{\exp(\tau\langle\mathbf{v}_{i},\mathbf{g}_{i}\rangle)}{\sum_{j=1}^{|\mathcal{B}|}\exp(\tau\langle\mathbf{v}_{i},\mathbf{g}_{j}\rangle)}$ (1)
@@
 3.3.2 Metrics
-[…]
+We evaluate the representation quality after image-transcriptomics alignment with Recall@$k$ for $k=\{1,5,10\}$, which measures the proportion of relevant items retrieved within the top $k$ retrieved samples.
+We report both image-to-gene (I2G) retrieval, where an image query retrieves its corresponding gene-expression profile, and gene-to-image (G2I) retrieval, where a gene-expression query retrieves the matching histological image.
 3.4 Downstream evaluations
@@
 The benchmark comprises nine tasks that span various tissue types and cancer subtypes from both 10x xenium and visium datasets.
-Each task predicts the expression of the top 50 most variable genes from 224$\times$224 pixel histology patches 112 $\mu$m $\times$ 112 $\mu$m at 20x magnification.
+Each task predicts the expression of the top 50 most variable genes from $\times224$ pixel histology patches $\mum ×\times 112 μ\mum$ at 20x magnification.
 The features extracted at the patch level are then passed to a ridge regression model for gene expression prediction.